# Beyond Numeric Awards: In-Context Dueling Bandits with LLM Agents

### v3

@@ B.3.1 Useful Assumptions and Lemmas for Dueling Bandits
With the TO property satisfied, we assume the preference matrix $P$ further satisfies the following two standard properties [Yue and Joachims, 2009, 2011, Yue et al., 2012].
@@
Note that the Bradley-Terry-Luce (BTL) model [Bradley and Terry, 1952] used in our experiments 3.1 satisfies Assumption 2 and 3. We restate the following theoretical guarantees for IF2 that is useful in the proof of Theorem 4.2. Let $\epsilon_{\mathrm{bad}}\coloneq\min_{b\neq b^{*}}\epsilon(b,b^{*})$.
Lemma 1 (Theorem 2 in [Yue et al., 2012]).
@@
The following expected regret bound achieved by IF2 is tight up to multiplicative constants, as indicated by the lower bound (Theorem 4) in [Yue et al., 2012] such that any algorithm Alg for DB satisfies $\mathbb{E}[\mathsf{SR}(\textsc{Alg})]=\Omega\left((K/\epsilon_{\mathrm{bad}})\log T\right)$.
@@ B.3.2 Theoretical Guarantees of LEAD
Vulnerability of Standalone LLM Agents. Inspired by the adversarial corruptions framework introduced in [Hajiesmaili et al., 2020] for the classic MAB problem, we investigate the vulnerability of standalone LLM agents in the DB setting under adversarial prompts. We consider an attacker with a budget $\Phi(T)$ who employs the following strategy: whenever the LLM agent selects the optimal arm $b^{*}$ for comparison, the attacker manipulates the input prompt to the LLM to eliminate $b^{*}$ from the duel with probability $p$ (where $0<p\leq 1$ is a constant), subject to the constraint of performing at most $\Phi(T)$ attacks over $T$ rounds. This adversarial strategy compels the LLM agent to select suboptimal arms, resulting in poor performance, as formalized in the following theorem with Assumption 4.
@@
Suppose at each step $t\leq T$, aligning with the design of IF2 in [Yue et al., 2012], $\hat{P}_{t}$ is estimated such that each $\smash{\hat{P}_{i,j}}$ is the fraction of number of comparisons when $b_{i}$ was the winner out of all previous $t$ comparisons. Define a confidence interval $\smash{\hat{C}_{t}\coloneq(\hat{P}_{t}-c_{t},\hat{P}_{t}+c_{t})}$ where $\smash{c_{t}\coloneq\sqrt{\log(1/\delta)/t}}$. Before proceeding to prove Theorem 4.2, we first state a useful lemma from [Yue et al., 2012] as a result of the Hoeffding’s inequality [Hoeffding, 1994].
Lemma 2 (Generalized Lemma 1 in [Yue et al., 2012]).
@@
Note that Lemma 2 can be directly implied by Lemma 1 in [Yue et al., 2012]. Now, under Assumption 2 and 3 such that the preference matrix $P$ satisfies the SST and STI properties, we prove Theorem 4.2.
@@
Weak Regret Bound. The first $T_{\mathrm{LLM}}$ steps induce accumulated weak regret of at most $O(T_{\mathrm{LLM}})$. According to [Yue et al., 2012], IF2 plays $O(K)$ matches (comparisons) in expectation. Thus, the expected number of rounds of calling IF2 Procedure is $O(\log T/\log(K\log T))$. Applying Lemma 2, with $\smash{O\left((1/\epsilon_{1,2}^{2})\log(K\log T)\right)}$ (by setting a hyper-parameter $\epsilon=\epsilon_{1,2}$) comparisons between two arms, since the best arm $b^{*}$ is always included in each comparison, the best arm $b^{*}$ is correctly identified with probability at least $1-1/(K\log T)^{2}$. This procedure leads to no weak regret since $b^{*}$ suggested by the LLM agent is always included as the incumbent arm in future comparisons.
@@ Appendix C Prompt Design and Supplementary Results
C.1 LLM Experimental Results
@@ C.1.1 Environments
In transitive instances, the preference matrices are constructed using the Bradley-Terry-Luce (BTL) model [Bradley and Terry, 1952, Yue et al., 2012], with a generalized form known as the Plackett-Luce model [Plackett, 1975]. In this model, each arm is associated with a utility parameter $\theta(i)>0$, where $i$ represents the rank of the arm (i.e., $\theta(1)$ corresponds to the best arm, $\theta(2)$ corresponds to the second best arm, and so on). For any pair of arms $b_{i}$ and $b_{j}$, the probability of $b_{i}$ being preferred over $b_{j}$ is determined by $P{(i\succ j)}=\theta(i)/(\theta(i)+\theta(j))$. Setting the number of arms $K=5$, we randomize the order of the arms to prevent selection bias, resulting in the following arm ordering: $b_{5}\succ b_{3}\succ b_{2}\succ b_{1}\succ b_{4}$. We use two instances: Transitive-Easy and Transitive-Hard, with their respective $\theta$ parameters given by:
@@
Note that the datasets generated in this way satisfy the Strong Stochastic Transitivity (SST) and Stochastic Triangle Inequality (STI) properties [Yue et al., 2012] (see Appendix B.3.1 for more details). The settings of the used BTL model also imply the existence of a Condorcet winner.
@@ C.1.2 Design of Prompts
C.1.3 Exemplars of GPT-4 Turbo and o1-preview
@@
C.2 Supplementary Experiments
C.2.1 Comparisons with Different Metrics
@@
C.2.2 Duel Selection Trajectory

### v4

@@ B.3.1 Useful Assumptions and Lemmas for Dueling Bandits
With the TO property satisfied, we assume the preference matrix $P$ further satisfies the following two standard properties (Yue and Joachims, 2009, 2011; Yue et al., 2012).
@@
Note that the Bradley-Terry-Luce (BTL) model (Bradley and Terry, 1952) used in our experiments C.1 satisfies Assumption 2 and 3. We restate the following theoretical guarantees for IF2 that is useful in the proof of Theorem 4.2. Let $\epsilon_{\mathrm{bad}}\coloneq\min_{b\neq b^{*}}\epsilon(b,b^{*})$.
Lemma 1 (Theorem 2 in (Yue et al., 2012)).
@@
The following expected regret bound achieved by IF2 is tight up to multiplicative constants, as indicated by the lower bound (Theorem 4) in (Yue et al., 2012) such that any algorithm Alg for DB satisfies $\mathbb{E}[\mathsf{SR}(\textsc{Alg})]=\Omega\left((K/\epsilon_{\mathrm{bad}})\log T\right)$.
@@ B.3.2 Theoretical Guarantees of LEAD
Vulnerability of standalone LLM agents. Inspired by the adversarial corruptions framework introduced in (Hajiesmaili et al., 2020) for the classic MAB problem, we investigate the vulnerability of standalone LLM agents in the DB setting under adversarial prompts. We consider an attacker with a budget $\Phi(T)$ who employs the following strategy: whenever the LLM agent selects the optimal arm $b^{*}$ for comparison, the attacker manipulates the input prompt to the LLM to eliminate $b^{*}$ from the duel with probability $p$ (where $0<p\leq 1$ is a constant), subject to the constraint of performing at most $\Phi(T)$ attacks over $T$ rounds. This adversarial strategy compels the LLM agent to select suboptimal arms, resulting in poor performance, as formalized in the following theorem with Assumption 4.
@@
Suppose at each step $t\leq T$, aligning with the design of IF2 in (Yue et al., 2012), $\hat{P}_{t}$ is estimated such that each $\smash{\hat{P}_{i,j}}$ is the fraction of number of comparisons when $b_{i}$ was the winner out of all previous $t$ comparisons. Define a confidence interval $\smash{\hat{C}_{t}\coloneq(\hat{P}_{t}-c_{t},\hat{P}_{t}+c_{t})}$ where $\smash{c_{t}\coloneq\sqrt{\log(1/\delta)/t}}$. Before proceeding to prove Theorem 4.2, we first state a useful lemma from (Yue et al., 2012) as a result of the Hoeffding’s inequality (Hoeffding, 1994).
Lemma 2 (Generalized Lemma 1 in (Yue et al., 2012)).
@@
Note that Lemma 2 can be directly implied by Lemma 1 in (Yue et al., 2012). Now, under Assumption 2 and 3 such that the preference matrix $P$ satisfies the SST and STI properties, we prove Theorem 4.2.
@@
Weak regret bound. The first $T_{\mathrm{LLM}}$ steps induce accumulated weak regret of at most $O(T_{\mathrm{LLM}})$. According to (Yue et al., 2012), IF2 plays $O(K)$ matches (comparisons) in expectation. Thus, the expected number of rounds of calling IF2 Procedure is $O(\log T/\log(K\log T))$. Applying Lemma 2, with $\smash{O\left((1/\epsilon_{1,2}^{2})\log(K\log T)\right)}$ (by setting a hyper-parameter $\epsilon=\epsilon_{1,2}$) comparisons between two arms, since the best arm $b^{*}$ is always included in each comparison, the best arm $b^{*}$ is correctly identified with probability at least $1-1/(K\log T)^{2}$. This procedure leads to no weak regret since $b^{*}$ suggested by the LLM agent is always included as the incumbent arm in future comparisons.
@@ Appendix C Prompt Design and Supplementary Results
C.1 Implementation Details of Experiments
Prompts and configurations of LLMs. We employ an interactive zero-shot chain-of-thought (CoT) prompt $\mathtt{Prompt}(P,H_{t},R)$, as defined in Section 2, which describes the problem setting $P$, externally summarized interaction history $H_{t}$ and reasoning instructions $R$. We adopt the prompting template and LLM configurations that lead to the best performance among all prompt variations explored in recent studies (Krishnamurthy et al., ; Nie et al., 2024) for MAB problem. The LLM agents interact with dueling bandit environments in a round-based manner, with the prompt guiding their decision-making process. We conduct experiments with five LLMs (Achiam et al., 2023; Touvron et al., 2023): GPT-3.5 Turbo, GPT-4, GPT-4 Turbo, Llama 3.1 (8B), and o1-preview. Our use of these scientific artifacts is in accordance with their respective licenses. Note that we skip the GPT-4o version which is primarily developed for multimodal tasks and has the same intelligence as GPT-4 Turbo. The detailed prompt is provided in Appendix C.2.2.
Baselines. We compare LLMs against nine well-established baseline algorithms to evaluate their efficacy. The baselines include Interleaved Filter (IF2(Yue et al., 2012), Beat the Mean (BTM(Yue and Joachims, 2011), Sensitivity Analysis of VAriables for Generic Exploration (SAVAGE(Urvoy et al., 2013), Relative Upper Confidence Bound (RUCB(Zoghi et al., 2014a), Relative Confidence Sampling (RCS(Zoghi et al., 2014b), Relative Minimum Empirical Divergence (RMED(Komiyama et al., 2015), Versatile Dueling Bandits (VDB(Saha and Gaillard, 2022), Self-Sparring (Sui et al., 2017), and Double Thompson Sampling (DTS(Wu and Liu, 2016). Each of these algorithms employs distinct strategies for selecting arms and estimating preferences, with the ultimate goal of efficiently identifying the Condorcet winner. We assess the performance of LLMs and baseline algorithms using strong regret and weak regret metrics defined in Section 2. We use $\gamma=0.5$ for BTM, $f(K)=0.3K^{1.01}$ for RMED, $\eta=1$ for Self-Sparring, and $\alpha=0.51$ for RUCB, RCS and DTS.
[…]
C.2 LLM Experimental Results
@@ C.2.1 Environments
In transitive instances, the preference matrices are constructed using the Bradley-Terry-Luce (BTL) model (Bradley and Terry, 1952; Yue et al., 2012), with a generalized form known as the Plackett-Luce model (Plackett, 1975). In this model, each arm is associated with a utility parameter $\theta(i)>0$, where $i$ represents the rank of the arm (i.e., $\theta(1)$ corresponds to the best arm, $\theta(2)$ corresponds to the second best arm, and so on). For any pair of arms $b_{i}$ and $b_{j}$, the probability of $b_{i}$ being preferred over $b_{j}$ is determined by $P{(i\succ j)}=\theta(i)/(\theta(i)+\theta(j))$. Setting the number of arms $K=5$, we randomize the order of the arms to prevent selection bias, resulting in the following arm ordering: $b_{5}\succ b_{3}\succ b_{2}\succ b_{1}\succ b_{4}$. We use two instances: Transitive-Easy and Transitive-Hard, with their respective $\theta$ parameters given by:
@@
Note that the datasets generated in this way satisfy the Strong Stochastic Transitivity (SST) and Stochastic Triangle Inequality (STI) properties (Yue et al., 2012) (see Appendix B.3.1 for more details). The settings of the used BTL model also imply the existence of a Condorcet winner.
@@ C.2.2 Design of Prompts
C.2.3 Exemplars of GPT-4 Turbo and o1-preview
@@
C.3 Supplementary Experiments
C.3.1 Comparisons with Different Metrics
@@
C.3.2 Duel Selection Trajectory